OpenStudy (fanduekisses):

I have a question: Is there a method to determine the symmetry of a function ALGEBRAICALLY or not graphically?

OpenStudy (fanduekisses):

For example: r = 3 cos 5θ

OpenStudy (kirbykirby):

you can check if \(f(\theta)=f(-\theta)\)

OpenStudy (fanduekisses):

I get r= - 3 cos5 θ

OpenStudy (fanduekisses):

the question asks me Determine if the graph is symmetric about the x-axis, the y-axis, or the origin.

OpenStudy (princeharryyy):

f(-theta)=f(theta)=> function is symmetric if, f(-theta)=-f(theta)=>function is retricemetric

OpenStudy (princeharryyy):

let the function be y=f(x), replce x by -x in the problem if f(-x) comes out to be f(x) yes the function is symmetric about y axis otherwise not

OpenStudy (anonymous):

To me, to polar equation like that, cos is always symmetric about x-axis, and sin is y-axis.

OpenStudy (fanduekisses):

Oh ok and how do I determine if it is symmetric about the x-axis or origin ?

OpenStudy (princeharryyy):

A graph will have symmetry about the x-axis if we get an equivalent equation when all the y’s are replaced with y.

OpenStudy (princeharryyy):

solvereplacing x and y's both for origin

OpenStudy (fanduekisses):

ohhhhhh ok thanks so much guys♥